The number density of electrons in a metal is 9.64×1028m−3 . What is the drift
velocity of electrons in a wire made up of this metal when a current of 4A flows in
it given that its cross-sectional area is 2.0 mm2? How long will it take for an
electron to travel a distance of 0.10 m in this wire?
1
Expert's answer
2019-03-22T10:36:34-0400

The current which occurs due to electron shift can be calculated as:


"I=nevS"

So, we can determine velocity:


"v=\frac{I}{neS}=\frac{4}{9.64*10^{28}*1.6*10^{-19}*2*10^{-6}}=130 m/s"

These electrons, will travel 0.1 m in:


"t=\frac{L}{v}=\frac{0.1}{130}=0.77 ms"